OpenStudy (anonymous):

Solve the following system. x^2 + y^2 = 13 2x - y = 4

OpenStudy (anonymous):

@Chlorophyll

OpenStudy (anonymous):

(3, 2) and (1/5, -3.6) ??

OpenStudy (ajprincess):

y=2x-4 x^2+(2x-4)^2=13 x^2+4x^2-16x+16=13 5x^2-16x+3=0 5x^2-15x-x+3=0 5x(x-3)-(x-3)=0 (5x-1)(x-3)=0 x=1/5 or x=3

OpenStudy (anonymous):

y = 2x - 4 => x^2 + ( 2x -4 ) ^2 = 13

OpenStudy (anonymous):

so what are the points? (?,?)(?,?)

OpenStudy (anonymous):

From @ajprincess's correct solution x = 3 => y = 2x - 4 = 2 => ( 3,2)

OpenStudy (ajprincess):

when x=3 y=2*3-4 =6-4 =2 when x=1/5 y=2*1/5-4 =2/5-4 =-18/5 =-3.6

OpenStudy (anonymous):

Similarly for x = 1/5 --> y = ...

OpenStudy (anonymous):

but there's two sets

OpenStudy (anonymous):

That's why you should look at airprincess's solution, will you!